# Nonlinear MPC for Feedback-Interconnected Systems: a Suboptimal and Reduced-Order Model Approach

### v1

@@ IV-B Reduced System Analysis
IV-C Proof of Theorem 1
The proof of Theorem 1 consists in merging the results achieved in Sections IV-A and IV-B to invoke the generic timescale separation result formalized in Theorem 2 (cf. Appendix VI-A). In order to apply this result, we need to check that its requirements are satisfied. First, by observing the requirement in (27), we need a Lyapunov function proving the exponential stability of the origin for the boundary-layer system (19) and we achieve it by Lemma 1. Second, by observing the requirement in (28), we need a Lyapunov function proving the exponential stability of $x_{\star}$ for our reduced system (21) and we achieve it by Assumption 4. Finally, we need the Lipschitz continuity of the equilibrium function $\operatorname{\mathrm{col}}(\xi_{\text{eq}}(x,\Pi z_{\star}(x)),z_{\star}(x))$ and the fast system dynamics as well as the conditions (23a) and (29) on the slow dynamics. These conditions are satisfied by Assumptions 12, and 3. Thus, we guarantee that $(x_{\star},\xi_{\text{eq}}(x_{\star},\Pi z_{\star}(x_{\star})),z_{\star}(x_{\star}))$ is a globally exponentially stable equilibrium of system (16) by invoking Theorem 2 and the proof concludes.
@@ V NUMERICAL SIMULATIONS
In this section, we present numerical simulations to validate the proposed Suboptimal and Reduced-Order MPC control strategy. In particular, we consider the actuated pendulum introduced in Example 1, whose physical parameters values are reported in Table I.
@@
$f_{R}(x,u,\delta):=\begin{bmatrix}x_{1}+\delta x_{2}\\ x_{2}+\delta\left(-\frac{g}{l}\sin(x_{1})-\frac{\beta}{ml^{2}}x_{2}+\frac{K_{t}}{ml^{2}}\frac{R}{K_{e}}u\right)\end{bmatrix}\!$.
We consider a finite-horizon optimal control problem (10) with cost functions given by
$\ell(x,u):=x^{\top}Qx+u^{\top}Ru,\quad\ell_{T}(x):=x^{\top}Q_{f}x,$
with $Q=Q_{f}=\operatorname{\mathrm{diag}}(100,0.1)$, $R=0.01$ and the prediction horizon $T=T_{s}\delta$ with $T_{s}=0.5\text{s}$. The input is subject to a saturation constraint $\mathcal{U}=\{u\in\mathbb{R}\mid|u|\leq 24\}$ and the MPC problem is addressed through a suboptimal scheme. We compare the proposed approach (Suboptimal and Red. Order), where the MPC optimizer relies on a reduced-order prediction model while interacting with the full-order plant, with two benchmark settings: Suboptimal and Full-Order and Optimal and Full-Order, where the MPC operates suboptimally and optimally, respectively. In both these benchmarks, the plant dynamics is reduced and, thus, coincides with the reduced-order prediction model, allowing the optimizer to fully capture the system behavior. These strategies are compared in three scenarios in which different $\delta$ are used. As predicted by Theorem 1, Fig. 5 shows that the closed-loop system exhibits exponential convergence towards the desired equilibrium configuration when a sufficiently small $\delta=0.01s$ is used. Fig. 6 highlights the importance of timescale separation. Indeed, for a larger value of $\delta$ ($\delta=0.1s$), although the suboptimal MPC optimizer operating on the full-order model still provides acceptable closed-loop behavior, the one operating on the reduced-order model exhibits a performance degradation. Finally, Fig. 7 further confirms the need for timescale separation. Indeed, in this case with $\delta=0.2s$, only the optimal MPC optimizer operating on the full-order model guarantees good closed-loop performance, while both the suboptimal MPC optimizers (operating on the full-order and reduced-order models) lead to a significant performance degradation.
Figure 5: Closed-loop trajectories of the target states $\theta$ and $\omega$ (left) and angular position error in semilog scale (right) for $\delta=0.01$ s.
Figure 6: Closed-loop trajectories of the target states $\theta$ and $\omega$ (left) and angular position error in semilog scale (right) for $\delta=0.1$ s.
Figure 7: Closed-loop trajectories of the target states $\theta$ and $\omega$ (left) and angular position error in semilog scale (right) for $\delta=0.2$ s.
@@ VI CONCLUSIONS
We proposed a suboptimal and reduced-order MPC scheme for discrete-time interconnected systems. The approach relies on a reduced-order model that omits part of the plant dynamics, enabling a lighter computational burden while maintaining formal closed-loop stability guarantees. In particular, we showed that closed-loop stability can be preserved by properly tuning a timescale parameter. Numerical simulations on an actuated pendulum confirmed the effectiveness of our approach.
APPENDIX
VI-A Global Exponential Stability of Two-Time-Scales Systems
The following result provides sufficient conditions for the global exponential stability of two-time-scale discrete-time systems. It is a slight extension of [22, Thm. II.5], since here we allow for generic slow dynamics in place of the specific form $x_{{{t}+1}}=x_{{t}}+\delta f(x_{{t}},w_{t})$ considered there.
Theorem 2
Consider the system
$x_{{{t}+1}}$ $=f(x_{{t}},w_{t},\delta)$ (22a)
$w_{{t}+1}$ $=G(w_{t},x_{{t}},\delta),$ (22b)
with $x_{{t}}\in\mathbf{X}\subseteq\mathbb{R}^{n}$, $w_{t}\in\boldsymbol{\Xi}\subseteq\mathbb{R}^{m}$, $f:\mathbf{X}\times\boldsymbol{\Xi}\times\mathbb{R}\rightarrow\mathbf{X}$, $G:\boldsymbol{\Xi}\times\mathbf{X}\times\mathbb{R}\rightarrow\boldsymbol{\Xi}$, $\delta>0$. Assume that there exist $L_{f},L_{g},\bar{\delta}_{1}>0$ such that for all $\delta\in(0,\bar{\delta}_{1})$, it holds
$\left\|f(x,w,\delta)\!-\!f(x^{\prime},w^{\prime},\delta)\right\|$ $\leq L_{f}\left\|x\!-\!x^{\prime}\right\|+\delta L_{f}\left\|w\!-\!w^{\prime}\right\|$ (23a)
$\left\|G(w,x,\delta)\!-\!G(w^{\prime},x^{\prime},\delta)\right\|$ $\!\leq\!L_{g}(\left\|w\!-\!w^{\prime}\right\|\!+\!\left\|x\!-\!x^{\prime}\right\|)\!\!$ (23b)
for all $x,x^{\prime}\in\mathbf{X}$, $w,w^{\prime}\in\boldsymbol{\Xi}$. Assume that there exists an $L_{h}$-Lipschitz continuous function $h:\mathbf{X}\to\boldsymbol{\Xi}$ such that
$h(x)=G(h(x),x,\delta),$ (24)
for all $x\!\in\!\mathbf{X}$. Assume also that there exists $x_{\star}\in\mathbf{X}$ such that $x_{\star}=f(x_{\star},h(x_{\star}),\delta)$, for all $\delta>0$. Then, let
$x_{{{t}+1}}=f(x_{{t}},h(x_{{t}}),\delta)$ (25)
be the reduced system and
$\psi_{{t}+1}=G(\psi_{t}+h(x),x,\delta)-h(x)$ (26)
be the boundary layer system with $\psi_{t}\in\boldsymbol{\Xi}$.
Assume that there exists a continuous function $U:\boldsymbol{\Xi}\to\mathbb{R}$ such that, for all $\delta\in(0,\bar{\delta}_{2})$, there exist $b_{1},b_{2},b_{3},b_{4}>0$ such that for all $\psi,\psi_{1},\psi_{2}\in\boldsymbol{\Xi}$, $x\in\mathbf{X}$, it holds
$b_{1}\left\|\psi\right\|^{2}\leq U(\psi)\leq b_{2}\left\|\psi\right\|^{2}$ (27a)
$U(G(\psi+h(x),x,\delta)-h(x))-U(\psi)\leq-b_{3}\left\|\psi\right\|^{2}$ (27b)
$U(\psi_{1})-U(\psi_{2})\leq b_{4}\left\|\psi_{1}-\psi_{2}\right\|(\left\|\psi_{1}\right\|+\left\|\psi_{2}\right\|).$ (27c)
Assume also that there exists a continuous function $W:\mathbf{X}\to\mathbb{R}$ and $\bar{\delta}_{3}>0$ such that, for all $\delta\in(0,\bar{\delta}_{3})$, it holds
$W(f(x,h(x),\delta))-W(x)\leq-\delta c_{3}\left\|x-x_{\star}\right\|^{2}$ (28b)
$W(x_{1})-W(x_{2})\leq c_{4}\left\|x_{1}-x_{2}\right\|(\left\|x_{1}-x_{\star}\right\|+\left\|x_{2}-x_{\star}\right\|),$ (28c)
for all $x,x_{1},x_{2},x_{3}\in\mathbf{X}$ and some $c_{1},c_{2},c_{3},c_{4}>0$. Finally, assume that, for all $x\!\in\!\mathbf{X}$ and $\delta\!>\!0$, it holds
$\left\|f(x,h(x),\delta)-x\right\|$ $\leq\delta L_{f}\left\|x-x_{\star}\right\|.$ (29)
Then, there exist $\bar{\delta}\!\in\!(0,\min\{\bar{\delta}_{1},\bar{\delta}_{2},\bar{\delta}_{3}\})$, $a_{1}>0$, and $a_{2}>0$ such that, for all $\delta\in(0,\bar{\delta})$, it holds
$\left\|\begin{bmatrix}x_{{t}}-x_{\star}\\ w_{t}-h(x_{{t}})\end{bmatrix}\right\|\leq a_{1}\left\|\begin{bmatrix}x^{0}-x_{\star}\\ w^{0}-h(x^{0})\end{bmatrix}\right\|e^{-a_{2}t},$
for all $(x^{0},w^{0})\in\mathbf{X}\times\boldsymbol{\Xi}$.
Proof:
Let us introduce $\tilde{w}_{t}:=w_{t}-h(x_{{t}})\!$ and accordingly rewrite system (22) as
$x_{{{t}+1}}$ $=f(x_{{t}},\tilde{w}_{t}+h(x_{{t}}),\delta)$ (30a)
$\tilde{w}_{{t}+1}$ $=G(\tilde{w}_{t}+h(x_{{t}}),x_{{t}},\delta)-h(x_{{{t}+1}}).$ (30b)
Select $W$ as in (28). By evaluating $\Delta W(x):=W(f(x,\tilde{w}+h(x),\delta))-W(x)$ along the trajectories of (30a), we obtain
$\Delta W(x)$ $=W(f(x,\tilde{w}+h(x),\delta))-W(x)$
$\stackrel{{\scriptstyle(a)}}{{=}}W(f(x,h(x),\delta))-W(x)$
$\hskip 11.38092pt+W(f(x,\tilde{w}+h(x),\delta))-W(f(x,h(x),\delta))$
$\stackrel{{\scriptstyle(b)}}{{\leq}}-\delta c_{3}\left\|x-x_{\star}\right\|^{2}+W(f(x,\tilde{w}+h(x),\delta))$
$\hskip 11.38092pt-W(f(x,h(x),\delta))$
$\stackrel{{\scriptstyle(c)}}{{\leq}}-\delta c_{3}\left\|x\!-\!x_{\star}\right\|^{2}\!+\!\delta c_{4}L_{f}\left\|\tilde{w}\right\|\!\left\|f(x,h(x),\delta)\!-\!x_{\star}\right\|$
$\hskip 11.38092pt+\delta c_{4}L_{f}\left\|\tilde{w}\right\|\left\|f(x,\tilde{w}+h(x),\delta)-x_{\star}\right\|,$ (31)
where in $(a)$ we add $\pm W(f(x,h(x),\delta))$, in $(b)$ we use (28b) to bound the first two terms, in $(c)$ we use (28c) and (23a). By using (23a), (29), and the triangle inequality, we note that
$\|f(x,\tilde{w}+h(x),\delta)-x_{\star}\|\leq(1\!+\!\delta L_{f})\left\|x-x_{\star}\right\|\!+\!\delta L_{f}\left\|\tilde{w}\right\|\!$. (32a)
$\left\|f(x,h(x),\delta)-x_{\star}\right\|\leq(1+\delta L_{f})\left\|x-x_{\star}\right\|.$ (32b)
Using inequalities (32a) and (32b) we then bound (31) as
$\Delta W(x)$ $\leq-\delta c_{3}\left\|x-x_{\star}\right\|^{2}+\delta^{2}k_{3}\left\|\tilde{w}\right\|^{2}$
$\hskip 11.38092pt+(\delta k_{1}+\delta^{2}k_{2})\left\|\tilde{w}\right\|\left\|x-x_{\star}\right\|,$ (33)
introducing $k_{1}:=2c_{4}L_{f},k_{2}:=2c_{4}L_{f}^{2},k_{3}:=c_{4}L_{f}^{2}$. Now, let us introduce $\Delta h:\mathbf{X}\times\boldsymbol{\Xi}\times\mathbb{R}\to\boldsymbol{\Xi}$ defined as
$\Delta h(x,\tilde{w},\delta):=-h(f(x,\tilde{w}+h(x),\delta))+h(x).$ (34)
[…]
$U(w~)\displaystyle\Delta$ U($\tilde${w})-U($\tilde${w})
[…]
$\hskip 11.38092pt-U(G(\tilde{w}+h(x),x,\delta)-h(x))$
$\hskip 11.38092pt+U(G(\tilde{w}+h(x),x,\delta)-h(f(x,\tilde{w}+h(x),\delta)))$
$\stackrel{{\scriptstyle(a)}}{{\leq}}-b_{3}\left\|\tilde{w}\right\|^{2}-U(G(\tilde{w}+h(x),x,\delta)-h(x))$
$\hskip 11.38092pt+U(G(\tilde{w}+h(x),x,\delta)-h(f(x,\tilde{w}+h(x),\delta)))$
$\stackrel{{\scriptstyle(b)}}{{\leq}}-b_{3}\left\|\tilde{w}\right\|^{2}+b_{4}\left\|\Delta h(x,\tilde{w},\delta)\right\|$
$\hskip 44.10185pt\times\left\|G(\tilde{w}+h(x),x,\delta)-h(f(x,\tilde{w}+h(x),\delta))\right\|$
$\hskip 11.38092pt+b_{4}\left\|\Delta h(x,\tilde{w},\delta)\right\|\left\|G(\tilde{w}+h(x),x,\delta)-h(x)\right\|$
$\stackrel{{\scriptstyle(c)}}{{\leq}}-b_{3}\left\|\tilde{w}\right\|^{2}+b_{4}\left\|\Delta h(x,\tilde{w},\delta)\right\|^{2}$
$\hskip 11.38092pt+2b_{4}\!\left\|\Delta h(x,\tilde{w},\delta)\right\|\!\left\|G(\tilde{w}+h(x),x,\delta)\!-\!h(x)\right\|\!,$ (35)
where in $(a)$ we exploit (27b) to bound the first two terms, in $(b)$ we use (27c) to bound the the difference of the last two terms, and in $(c)$ we use the triangle inequality. By definition of $\Delta h$ (cf. (34)) and Lipschitz continuity of $h$, we write
$\left\|\Delta h(x,\tilde{w},\delta)\right\|$ $\leq L_{h}\left\|f(x,\tilde{w}+h(x),\delta)-x\right\|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}L_{h}\left\|f(x,h(x),\delta)-x\right\|$
$\hskip 11.38092pt+L_{h}\left\|f(x,\tilde{w}+h(x),\delta)-f(x,h(x),\delta)\right\|$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\delta L_{h}L_{f}\left\|x-x_{\star}\right\|+\delta L_{h}L_{f}\left\|\tilde{w}\right\|$, (36)
where in $(a)$ we add and subtract $f(x,h(x),\delta)$ in the norm and apply the triangle inequality, while in $(b)$ we apply (29) to bound the first term and (23a) to bound the second one. Moreover, since $G(h(x),x,\delta)=h(x)$ (cf. (24)), we obtain
$\left\|G(\tilde{w}+h(x),x,\delta)-h(x)\right\|$
$=\left\|G(\tilde{w}+h(x),x,\delta)-G(h(x),x,\delta)\right\|\leq L_{G}\left\|\tilde{w}\right\|\!$, (37)
where the inequality is due to the Lipschitz continuity of $G$ (cf. (23b)). Using inequalities (36) and (37), we then bound (35) as
[…]
$\hskip 11.38092pt+2\delta b_{4}L_{h}L_{G}L_{f}\left\|x-x_{\star}\right\|\left\|\tilde{w}\right\|$
$\hskip 11.38092pt+2\delta^{2}b_{4}L_{h}^{2}L_{f}^{2}\left\|x-x_{\star}\right\|\left\|\tilde{w}\right\|+\delta^{2}b_{4}L_{h}^{2}L_{f}^{2}\left\|\tilde{w}\right\|^{2}$
$\hskip 11.38092pt+\delta^{2}b_{4}L_{h}^{2}L_{f}^{2}\left\|x-x_{\star}\right\|^{2}$
$\leq(-b_{3}+\delta k_{6}+\delta^{2}k_{7})\left\|\tilde{w}\right\|^{2}+\delta^{2}k_{8}\left\|x-x_{\star}\right\|^{2}$
$\hskip 11.38092pt+(\delta k_{4}+\delta^{2}k_{5})\left\|x-x_{\star}\right\|\left\|\tilde{w}\right\|$, (38)
where we introduce the constants
$k_{4}$ $:=2b_{4}L_{h}L_{G}L_{f},\hskip 1.42271pt$ $k_{5}:=2b_{4}L_{h}^{2}L_{f}^{2},\hskip 2.84544ptk_{6}:=2b_{4}L_{h}L_{G}L_{f},$
$k_{7}:=b_{4}L_{h}^{2}L_{f}^{2},\hskip 1.42271pt$ $k_{8}:=b_{4}L_{h}^{2}L_{f}^{2}.$
We pick the following Lyapunov candidate $V:\mathbf{X}\times\boldsymbol{\Xi}\to\mathbb{R}$:
[…]
where we define the matrix $Q(\delta)=Q(\delta)^{\top}\in\mathbb{R}^{2}$ as
$Q(\delta):=\begin{bmatrix}\delta c_{3}-\delta^{2}k_{8}&q_{21}(\delta)\\ q_{21}(\delta)&b_{3}-\delta k_{6}-\delta^{2}(k_{3}+k_{7})\end{bmatrix}$,
with $q_{21}(\delta):=-\frac{1}{2}(\delta(k_{1}+k_{4})+\delta^{2}(k_{2}+k_{5}))$. By Sylvester criterion, we know that $Q\succ 0$ if and only if
$c_{3}b_{3}>p(\delta),$ (40)
where the polynomial $p(\delta)$ is defined as
$p(\delta)$ $:=q_{21}(\delta)^{2}+\delta^{2}c_{3}k_{6}+\delta^{2}(\delta c_{3}(k_{3}+k_{7})+b_{3}k_{8})$
$\hskip 14.22636pt-\delta^{3}k_{6}k_{8}-\delta^{4}k_{8}(k_{3}+k_{7}).$ (41)
Since $p$ is a continuous function of $\delta$ and $\lim_{\delta\to 0}p(\delta)=0$, there exists $\bar{\delta}\in(0,\min\{\bar{\delta}_{1},\bar{\delta}_{2},\bar{\delta}_{3}\})$ so that (40) is satisfied for any $\delta\in(0,\bar{\delta})$. Under this choice of $\delta$ and denoting by $q>0$ the smallest eigenvalue of $Q(\delta)$, we bound (39) as
$\Delta V(x,\tilde{w})\leq-q\left\|x-x_{\star}\right\|^{2}+\left\|\tilde{w}\right\|^{2},$
which allows us to conclude, in view of [23, Theorem 13.2], that $(x_{\star},0)$ is an exponentially stable equilibrium point for system (30). The conclusion of the theorem follows by considering the definition of exponentially stable equilibrium point and by reverting to the original coordinates $(x,w)$. ∎
VI-B Proof of Lemma 1
The proof consists in combining the exponential stability properties of the fast dynamics (1b) stated in Assumption 2 with those of the optimizer dynamics (11) stated in Assumption 3, and in suitably handling the drift of the fast plant equilibrium $\xi_{\text{eq}}(x,\Pi(\tilde{z}_{t}+z_{\star}(x)))$ due to the variations of $\tilde{z}_{t}$. Purposely, let the function $\mathcal{U}:\boldsymbol{\Xi}\times\mathbf{Z}\to\mathbb{R}$ be defined as
[…]
where $U$ and $\mathcal{L}$ are introduced in (3) (cf. Assumption 2) and (14) (cf. Assumption 3), respectively, while $\kappa\!>\!0$ is a constant that we will set later. We note that (20a) and (20c) directly follow by combining (3a) with (14a) for the former and (3c) with (14c) for the latter. To prove (20b), we introduce a more compact notation by defining the shorthands
[…]
$\tilde{z}_{+}$ $:=\tilde{\mathcal{T}}(\tilde{z},x),\quad\tilde{\xi}_{+}^{\text{nom}}:=\tilde{g}(\tilde{\xi},x,\Pi(\tilde{z}+z_{\star}(x)),\delta).$ (43b)
[…]
$\stackrel{{\scriptstyle(a)}}{{\leq}}U(\tilde{\xi}_{+})-U(\tilde{\xi})-\kappa b_{3}\left\|\tilde{z}\right\|^{2}$
$\stackrel{{\scriptstyle(b)}}{{=}}U(\tilde{\xi}_{+}^{\text{nom}})\!-\!U(\tilde{\xi})\!-\!\kappa b_{3}\left\|\tilde{z}\right\|^{2}\!+\!U(\tilde{\xi}_{+})-U(\tilde{\xi}_{+}^{\text{nom}})$
$\stackrel{{\scriptstyle(c)}}{{\leq}}-a_{3}\left\|\tilde{\xi}\right\|^{2}\!-\!\kappa b_{3}\left\|\tilde{z}\right\|^{2}\!+\!U(\tilde{\xi}_{+})\!-\!U(\tilde{\xi}_{+}^{\text{nom}}),$ (44)
where in $(a)$ we apply (14b) (cf. Assumption 3), in $(b)$ we add $\pm U(\tilde{\xi}_{+}^{\text{nom}})$, while in $(c)$ we apply (3b). Now, we focus on the term $U(\tilde{\xi}_{+})-U(\tilde{\xi}_{+}^{\text{nom}})$. By using the bound (3c) and the definitions of $\tilde{\xi}_{+}$ (cf. (43a)) and $\tilde{\xi}_{+}^{\text{nom}}$ (cf. (43b)), we get
$U(\tilde{\xi}_{+})-U(\tilde{\xi}_{+}^{\text{nom}})$
[…]
$\hskip 11.38092pt+a_{4}\left\|\Delta\xi_{\text{eq}}(x,\tilde{z}_{+},\tilde{z})\right\|\left\|\tilde{g}(\tilde{\xi},x,\Pi(\tilde{z}+z_{\star}(x)),\delta)\right\|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}a_{4}\left\|\Delta\xi_{\text{eq}}(x,\tilde{z}_{+},\tilde{z})\right\|^{2}$
$\hskip 11.38092pt+2a_{4}\left\|\Delta\xi_{\text{eq}}(x,\tilde{z}_{+},\tilde{z})\right\|\left\|\tilde{g}(\tilde{\xi},x,\Pi(\tilde{z}+z_{\star}(x)),\delta)\right\|\!,\!\!$ (45)
where $(a)$ follows by the triangular inequality. By observing the definition of $\tilde{g}$ (cf. (18a)), we note that $\tilde{g}(0,x,\Pi(z_{\star}(x)))=0$ for all $x\in\mathbf{X}$ in light of (2) (cf. Assumption 2). Thus, by combining this property with the Lipschitz continuity of $\tilde{g}$ (cf. Assumption 1), it holds
$\left\|\tilde{g}(\tilde{\xi},x,\Pi(\tilde{z}+z_{\star}(x)),\delta)\right\|\leq L_{g}\left\|\tilde{\xi}\right\|+\left\|\tilde{z}\right\|$. (46)
By observing the definition of $\Delta\xi_{\text{eq}}$ (cf. (18c)) and using the Lipschitz continuity of $\xi_{\text{eq}}$ (cf. Assumption 2), we get
[…]
$\stackrel{{\scriptstyle(a)}}{{\leq}}L_{\xi}\left\|\tilde{\mathcal{T}}(\tilde{z},x)\right\|+L_{\xi}\left\|\tilde{z}\right\|$
$\stackrel{{\scriptstyle(b)}}{{\leq}}L_{\xi}(L_{\mathcal{T}}+1)\left\|\tilde{z}\right\|$, (47)
where in $(a)$ we use the Cauchy-Schwarz inequality, the fact that $\left\|\Pi\right\|=1$, the definition of $\tilde{z}_{+}$ (cf. (43b)), and the triangle inequality, while in $(b)$ we use the fact that $\tilde{\mathcal{T}}(0,x)=0$ for all $x\in\mathbf{X}$ (see (13) in Assumption 3) and the Lipschitz continuity of $\tilde{\mathcal{T}}$ (cf. Assumption 3). Therefore, by using (46) and (47), we further bound the right-hand side of (45) as
$U(\tilde{\xi}_{+})-U(\tilde{\xi}_{+}^{\text{nom}})\leq 2k_{1}\left\|\tilde{\xi}\right\|\left\|\tilde{z}\right\|+k_{2}\left\|\tilde{z}\right\|^{2},$ (48)
in which we introduce the constants $k_{1}:=a_{4}L_{\xi}L_{g}(L_{\mathcal{T}}+1)$ and $k_{2}:=a_{4}\left(2L_{\xi}(L_{\mathcal{T}}+1)+(L_{\xi}(L_{\mathcal{T}}+1))^{2}\right)$. In turn, by using (48), we bound the right-hand side of (44) as
[…]
$\stackrel{{\scriptstyle(a)}}{{=}}-\begin{bmatrix}\left\|\tilde{\xi}\right\|\\ \left\|\tilde{z}\right\|\end{bmatrix}^{\top}\underbrace{\begin{bmatrix}a_{3}&-k_{1}\\ -k_{1}&\kappa b_{3}-k_{2}\end{bmatrix}}_{H(\kappa)}\begin{bmatrix}\left\|\tilde{\xi}\right\|\\ \left\|\tilde{z}\right\|\end{bmatrix}$, (49)
where in $(a)$ we adopt a compact matrix notation. By Sylvester criterion, we have
$H(\kappa)\succ 0\iff a_{3}(\kappa b_{3}-k_{2})>k_{1}^{2}\iff\kappa>\frac{k_{1}^{2}}{a_{3}b_{3}}+\frac{k_{2}}{b_{3}}.$
The proof follows by accordingly setting $\kappa$ and denoting with $d_{3}>0$ the smallest eigenvalue of $H(\kappa)$.

### v2

@@ IV-B Reduced System Analysis
IV-C Sketch of proof of Theorem 1
The proof relies on timescale-separation arguments applied to the closed-loop system (17) to establish the stability of its equilibrium. More precisely, the analysis combines: (i) the stability properties of the boundary-layer system (cf. Lemma 1), (ii) those of the reduced system (cf. Assumption 4), and (iii) a careful treatment of the actual interconnected dynamics, rather than the “ideal” dynamics considered in these two auxiliary systems. Indeed, we recall that the boundary-layer system neglects the drift of its equilibrium, while, analogously, the reduced system assumes that the fast state lies at all times on the equilibrium manifold. Therefore, these two ingredients must be combined appropriately, together with a suitable tuning of $\delta$ and appropriate regularity assumptions on the dynamics and on the equilibrium map, in order to control the variation of $x_{{t}}$ and derive an upper bound $\bar{\delta}$ below which stability of the interconnected system is preserved.
@@ V NUMERICAL SIMULATIONS
In this section, we present numerical simulations to validate the proposed Suboptimal and Reduced-Order MPC control strategy. In particular, we consider the actuated pendulum introduced in Example 1, whose physical parameter values are reported in Table I. All simulations were conducted using the CasADi tool for numerical optimization [23].
@@
$f_{R}(x,u,\delta)\!:=\!\begin{bmatrix}x_{1}+\delta x_{2}\\ x_{2}\!+\!\!\delta\left(-\frac{g}{l}\sin(x_{1})\!-\!\frac{\beta R+K_{t}K_{e}}{ml^{2}R}x_{2}+\frac{K_{t}}{ml^{2}R}u\right)\end{bmatrix}\!$.
In all simulations, we consider $T=4$ and cost functions and saturation constraints given by
$\ell(x,u):=x^{\top}Qx+0.01u^{2},\quad\ell_{T}(x):=x^{\top}Q_{f}x$
with $Q=Q_{f}:=\operatorname{\mathrm{diag}}(100,0.1)$, $R:=0.01$. The system is subject to the actuation constraints: $u\in\mathcal{U}:=\{u\in\mathbb{R}\mid|u|\leq 12\}$. The MPC problem is addressed suboptimally through a single step of a projected gradient method.
We compare the proposed approach, where the suboptimal MPC optimizer relies on a reduced-order prediction model while interacting with the actual full-order plant, against two benchmark settings. In these benchmarks, the MPC operates either optimally or suboptimally, but the plant is artificially forced to operate under an ideal steady-state regime for its fast dynamics, i.e., $\xi_{{t}}=\xi_{\text{eq}}(x_{{t}},u_{{t}})$ for all $t\in\mathbb{N}$. Because this idealized plant behavior perfectly aligns with the reduced-order prediction model, no structural plant-model mismatch exists for the optimizer in these benchmarks. To accurately reflect this condition, we refer to these benchmarks as Optimal (Ideal Plant) and Suboptimal (Ideal Plant). These strategies are compared in three scenarios in which different $\delta$ are used. As predicted by Theorem 1, Fig. 5 shows that the closed-loop system exhibits exponential convergence towards the desired equilibrium configuration when a sufficiently small $\delta=0.01s$ is used. Fig. 6 highlights the importance of timescale separation. Indeed, for a larger value of $\delta$ ($\delta=0.12s$), the suboptimal MPC still provides acceptable closed-loop behavior when controlling the idealized matched plant but its performance degrades significantly when interacting with the actual full-order plant.
Finally, Fig. 7 further confirms the need for timescale separation. In this case, in which $\delta=0.2s$, only the exact optimal MPC applied to the idealized matched plant guarantees good closed-loop performance. Conversely, the suboptimal numerical optimization leads to a severe performance degradation even in the absence of plant-model mismatch, a deterioration that is further exacerbated when controlling the full-order plant.
Figure 5: Closed-loop trajectories of the target states $\theta$ and $\omega$ (left) and state error in semilog scale (right) for $\delta=0.01$ s.
Figure 6: Closed-loop trajectories of the target states $\theta$ and $\omega$ (left) and state error in semilog scale (right) for $\delta=0.12$ s.
Figure 7: Closed-loop trajectories of the target states $\theta$ and $\omega$ (left) and state error in semilog scale (right) for $\delta=0.2$ s.
@@ VI CONCLUSIONS
We proposed a suboptimal and reduced-order MPC scheme for discrete-time interconnected systems. We relied on a reduced-order model that omits part of the plant dynamics, enabling a lighter computational burden while maintaining closed-loop stability guarantees. In detail, we showed that stability is preserved by properly tuning a timescale parameter. Numerical tests on an actuated pendulum confirmed the effectiveness of our approach. A rigorous treatment of recursive feasibility is left for future work.